OpenStudy (anonymous):

differentiate: y=(2x^3-4x+1)^2/3

OpenStudy (campbell_st):

do you know how to use the chain rule..?

OpenStudy (anonymous):

no

OpenStudy (campbell_st):

ok... that make it very hard... because thats what you need for this question

OpenStudy (anonymous):

haay :/ how to use that ?

OpenStudy (campbell_st):

the chain rule is \[\frac{dy}{dx} = \frac{dy}{du} \times \frac{du}{dx}\] you would need to let \[u = 2x^3 -4x + 1\] and \[y = u^{\frac{2}{3}}\]

OpenStudy (campbell_st):

can you find du/dx...? and dy/du..?

OpenStudy (anonymous):

 is that ?